OpenStudy (anonymous):

Find a Cartesian equation for the curve and identify it. r=8tanxsecx

terenzreignz (terenzreignz):

tanxsecx? Are you sure it isn't \[\Large r = 8\tan(\theta) \sec(\theta)\]?

OpenStudy (anonymous):

yeah its r = 8 tan θ sec θ :d

terenzreignz (terenzreignz):

Okay, these are all you really need for that: \[\Large r= \sqrt{x^2+y^2}\]\[\Large \tan(\theta) = \frac{y}x\] \[\Large x = r\cos(\theta)\]

terenzreignz (terenzreignz):

scratch that, for this particular question, you don't actually need the first one \[\large r = \sqrt{x^2+y^2}\]

OpenStudy (anonymous):

(rcosθ)^2=8(rsinθ) is that right:d

terenzreignz (terenzreignz):

Actually, yes. But that equation is still not Cartesian, though I have a feeling you already know how to fix that ;)

terenzreignz (terenzreignz):

And also (as you well know, I guess) \[\Large y = r\sin(\theta)\]

OpenStudy (anonymous):

thanks.d

OpenStudy (anonymous):

x^2 = 8y right :d

terenzreignz (terenzreignz):

Yup. That's a parabola XD